Two vectors are given: π‘₯βƒ— = 3𝑖⃗ + 2𝑗⃗, 𝑦⃗ = 2𝑖⃗ βˆ’ 8𝑗⃗.


a) Draw these vectors in Cartesian coordinate system.


b) Represent both of these vectors in magnitude, direction notation.


c) Represent vector 2π‘₯βƒ— + 3𝑦⃗ in magnitude, direction notation.

1
Expert's answer
2021-11-14T17:10:49-0500

a) The diagram:



b) The magnitude-direction rotation:


"|x|=\sqrt{3^2+2^2}=3.6,\\\space\\
\theta_x=\arctan\frac{2}{3}=34°,\\\space\\
|y|=\sqrt{2^2+(-8)^2}=8.2,\\\space\\
\theta_y=\arctan\frac{-8}{2}=-76°."

c) The magnitude-direction rotation:


"|2x+3y|=\sqrt{(2·3+3·2)^2+(2·2+3(-8))^2}=23,\\\space\\
\theta_{x+y}=\arctan\frac{2·3+3·2}{2·3+3·2}=-59°."